# E2 TTS: Embarrassingly Easy Fully Non-Autoregressive Zero-Shot TTS

### v1

@@ 1 Introduction
Meanwhile, several fully NAR zero-shot TTS models have been proposed with promising results. Unlike AR-based models, fully NAR models enjoy fast inference based on parallel processing. NaturalSpeech 2 [14] and NaturalSpeech 3 [15] estimate the latent vectors of a neural audio codec based on diffusion models [16, 17]. Voicebox [18] and MatchaTTS [19] used a flow-matching model [20] conditioned by an input text. However, one notable challenge for such NAR models is how to obtain the alignment between the input text and the output audio, whose length is significantly different. NaturalSpeech 2, NaturalSpeech 3, and Voicebox used a frame-wise phoneme alignment for training the model. MatchaTTS, on the other hand, used monotonic alignment search (MAS) [21, 3] to automatically find the alignment between input and output. While MAS could alleviate the necessity of the frame-wise phoneme aligner, it still requires an independent duration model to estimate the duration of each phoneme during inference. More recently, E3 TTS [22] proposed using cross-attention from the input sequence, which required a carefully designed U-Net architecture [23]. As such, fully NAR zero-shot TTS models require either an explicit duration model or a carefully designed architecture. One of our findings in this paper is that such techniques are not necessary to achieve high-quality zero-shot TTS, and they are sometimes even harmful to naturalness. [^1] Concurrent with our work, Seed-TTS [24] proposed a diffusion model-based zero-shot TTS, named Seed-TTSDiT. Although it appears to share many similarities with our approach, the authors did not elaborate on the details of their model, making it challenging to compare with our work.
@@ 2.1 Training
Fig. 1 (a) provides an overview of E2 TTS training. Suppose we have a training audio sample $s$ with transcription $y=(c_{1},c_{2},...,c_{M})$, where $c_{i}$ represents the $i$-th character of the transcription. [^2] Alternatively, we can represent $y$ as a sequence of Unicode bytes [25]. First, we extract its mel-filterbank features $\hat{s}\in\mathbb{R}^{D\times T}$, where $D$ denotes the feature dimension and $T$ represents the sequence length. We then create an extended character sequence $\tilde{y}$, where a special filler token $\langle F\rangle$ is appended to $y$ to make the length of $\tilde{y}$ equal to $T$ [^3] We assume $M\leq T$, which is almost always valid.
@@ 2.3 Flow-matching-based mel spectrogram generator
where $p_{t}$ is the probability path at time $t$, $u_{t}$ is the designated vector field for $p_{t}$, $x_{1}$ symbolizes the random variable corresponding to the training data, and $q$ is the distribution of the training data. In the training phase, we construct both a probability path and a vector field from the training data, utilizing an optimal transport path: $p_{t}(x|x_{1})=\mathcal{N}(x|tx_{1},(1-(1-\sigma_{\rm min})t)^{2}I)$ and $u_{t}(x|x_{1})=(x_{1}-(1-\sigma_{\rm min})x)/(1-(1-\sigma_{\rm min})t)$. For inference, we apply an ordinary differential equation solver [27] to generate the log mel-filterbank features starting from the initial distribution $p_{0}$.
We adopt the same model architecture with the audio model of Voicebox (Fig. 2 of [18]) except that the frame-wise phoneme sequence is replaced into $\hat{y}$. Specifically, Transformer with U-Net style skip connection [18] is used as a backbone. The input to the mel spectrogram generator is $m\odot\hat{s}$, $\hat{y}$, the flow step $t$, and noisy speech $s_{t}$. $\hat{y}$ is first converted to character embedding sequence $\tilde{y}\in\mathbb{R}^{E\times T}$. Then, $m\odot\hat{s}$, $s_{t}$, $\tilde{y}$ are all stacked to form a tensor with a shape of $(2\cdot D+E)\times T$, followed by a linear layer to output a tensor with a shape of $D\times T$. Finally, an embedding representation, $\hat{t}\in\mathbb{R}^{D}$, of $t$ is appended to form the input tensor with a shape of $\mathbb{R}^{D\times(T+1)}$ to the Transformer. The Transformer is trained to output a vector field $v_{t}$ with the conditional flow-matching objective $\mathcal{L}^{\rm CFM}$.
@@ 2.5.1 Extension 1: Eliminating the need for transcription of audio prompts in inference
In certain application contexts, obtaining the transcription of the audio prompt can be challenging. To eliminate the requirement of the transcription of the audio prompt during inference, we introduce an extension, illustrated in Fig. 2. This extension, referred to as E2 TTS X1, assumes that we have access to the transcription of the masked region of the audio, which we use for $y$. During inference, the extended character sequence $\tilde{y}^{\prime}$ is formed without $y^{\rm aud}$, namely,
$\tilde{y}^{\prime}=(c^{\prime\prime}_{1},c^{\prime\prime}_{2},\ldots,c^{\prime\prime}_{M^{\rm text}},\underbrace{\langle F\rangle,\ldots,\langle F\rangle}_{\mathcal{T}\text{ times}}).$ (4)
@@
The transcription of the masked region of the training audio can be obtained in several ways. One method is to simply apply ASR to the masked region during training, which is straightforward but costly. In our experiment, we employed the Montreal Forced Aligner [28] to determine the start and end times of words within each training data sample. The masked region was determined in such a way that we ensured not to cut the word in the middle.
@@ 3.1 Training data
We utilized the Libriheavy dataset [30] to train our models. The Libriheavy dataset comprises 50,000 hours of read English speech from 6,736 speakers, accompanied by transcriptions that preserve case and punctuation marks. It is derived from the Librilight [31] dataset contains 60,000 hours of read English speech from over 7,000 speakers. For E2 TTS training, we used the case and punctuated transcription without any pre-processing. We also used a proprietary 200,000 hours of training data to investigate the scalability of E2 TTS model.
@@ 3.2 Model configurations
We constructed our proposed E2 TTS models using a Transformer architecture. The architecture incorporated U-Net [23] style skip connections, 24 layers, 16 attention heads, an embedding dimension of 1024, a linear layer dimension of 4096, and a dropout rate of 0.1. The character embedding vocabulary size was 399.555We used all characters and symbols that we found in the training data without filtering. The total number of parameters amounted to 335 million. We modeled the 100-dimensional log mel-filterbank features, extracted every 10.7 milliseconds from audio samples with a 24 kHz sampling rate. A BigVGAN [32]-based vocoder was employed to convert the log mel-filterbank features into waveforms. The masking length was randomly determined to be between 70% and 100% of the log mel-filterbank feature length during training. In addition, we randomly dropped all the conditioning information with a 20% probability for classifier-free guidance (CFG) [33]. All models were trained for 800,000 mini-batch updates with an effective mini-batch size of 307,200 audio frames. We utilized a linear decay learning rate schedule with a peak learning rate of $7.5\times 10^{-5}$ and incorporated a warm-up phase for the initial 20,000 updates. We discarded the training samples that exceeded 4,000 frames.
@@
In addition, we trained a regression-based duration model by following that of Voicebox [18]. It is based on a Transformer architecture, consisted of 8 layers, 8 attention heads, an embedding dimension of 512, a linear layer dimension of 2048, and a dropout rate of 0.1. The training process involved 75,000 mini-batch updates with 120,000 frames. We used this duration model for estimating the target duration for the fair comparison to the Voicebox baseline. Note that we will also show that E2 TTS is robust for different target duration in Section 3.6.
@@ 3.3 Evaluation data and metrics
In order to assess our models, we utilized the test-clean subset of the LibriSpeech-PC dataset [35], which is an extension of LibriSpeech [36] that includes additional punctuation marks and casing. We specifically filtered the samples to retain only those with a duration between 4 and 10 seconds. Since LibriSpeech-PC lacks some of the utterances from LibriSpeech, the total number of samples was reduced to 1,132, sourced from 39 speakers. For the audio prompt, we extracted the last three seconds from a randomly sampled speech file from the same speaker for each evaluation sample.
@@
Table 1: Objective results for LibriSpeech-PC test-clean evaluation set. WER is expressed in percentage. Our reproduction. LL, LH, and PP stand for Libriheavy, Librilight, and Proprietary, respectively.
@@
Table 3: Comparison between the basic E2 TTS and E2 TTS-E1. While the basic E2 TTS requires the transcription of the audio prompt during inference, E2 TTS X1 does not require it. WER is expressed in percentage.
@@ 3.4 Main results
In our experiments, we conducted a comparison between our E2 TTS models and four other models: Voicebox [18], VALL-E [8], and NaturalSpeech 3 [15]. We utilized our own reimplementation of the Voicebox model, which was based on the same model configuration with E2 TTS except that the Vicebox model is trained with frame-wise phoneme alignment. During the inference, we used CFG with a guidance strength of 1.0 for both E2 TTS and Voicebox. We employed the midpoint ODE solver with a number of function evaluations of 32.
@@
Table 4: The WER (%) and SIM-o of E2 TTS-E2 where a word is randomly replaced with the phoneme sequence during inference. Even when we replaced 50% of words into phoneme sequences, E2 TTS-E2 worked reasonably well. This indicates that we can specify the pronunciation of a new term without retraining.
@@ 3.5.2 Evaluation of the extension 2
In this experiment, we trained the E2 TTS X2 models with a phoneme replacement rate of 15%. During inference, we randomly replaced words in the test-clean dataset with phoneme sequences, with a probability ranging from 0% up to 50%.

### v2

@@ 1 Introduction
Meanwhile, several fully NAR zero-shot TTS models have been proposed with promising results. Unlike AR-based models, fully NAR models enjoy fast inference based on parallel processing. NaturalSpeech 2 [14] and NaturalSpeech 3 [15] estimate the latent vectors of a neural audio codec based on diffusion models [16, 17]. Voicebox [18] and Matcha-TTS [19] used a flow-matching model [20] conditioned by an input text. However, one notable challenge for such NAR models is how to obtain the alignment between the input text and the output audio, whose length is significantly different. NaturalSpeech 2, NaturalSpeech 3, and Voicebox used a frame-wise phoneme alignment for training the model. Matcha-TTS, on the other hand, used monotonic alignment search (MAS) [21, 3] to automatically find the alignment between input and output. While MAS could alleviate the necessity of the frame-wise phoneme aligner, it still requires an independent duration model to estimate the duration of each phoneme during inference. More recently, E3 TTS [22] proposed using cross-attention from the input sequence, which required a carefully designed U-Net architecture [23]. As such, fully NAR zero-shot TTS models require either an explicit duration model or a carefully designed architecture. One of our findings in this paper is that such techniques are not necessary to achieve high-quality zero-shot TTS, and they are sometimes even harmful to naturalness. [^1] Concurrent with our work, Seed-TTS [24] proposed a diffusion model-based zero-shot TTS, named Seed-TTSDiT. Although it appears to share many similarities with our approach, the authors did not elaborate on the details of their model, making it challenging to compare with our work.
@@ 2.1 Training
Fig. 1 (a) provides an overview of E2 TTS training. Suppose we have a training audio sample $s$ with transcription $y=(c_{1},c_{2},...,c_{M})$, where $c_{i}$ represents the $i$-th character of the transcription. [^2] Alternatively, we can represent $y$ as a sequence of Unicode bytes [25]. First, we extract its mel-filterbank features $\hat{s}\in\mathbb{R}^{D\times T}$, where $D$ denotes the feature dimension and $T$ represents the sequence length. We then create an extended character sequence $\hat{y}$, where a special filler token $\langle F\rangle$ is appended to $y$ to make the length of $\hat{y}$ equal to $T$ [^3] We assume $M\leq T$, which is almost always valid.
@@ 2.3 Flow-matching-based mel spectrogram generator
where $p_{t}$ is the probability path at time $t$, $u_{t}$ is the designated vector field for $p_{t}$, $x_{1}$ symbolizes the random variable corresponding to the training data, and $q$ is the distribution of the training data. In the training phase, we construct both a probability path and a vector field from the training data, utilizing an optimal transport path: $p_{t}(x|x_{1})=\mathcal{N}(x|tx_{1},(1-(1-\sigma_{\rm min})t)^{2}I)$ and $u_{t}(x|x_{1})=(x_{1}-(1-\sigma_{\rm min})x)/(1-(1-\sigma_{\rm min})t)$. For inference, we apply an ordinary differential equation (ODE) solver [27] to generate the log mel-filterbank features starting from the initial distribution $p_{0}$.
We adopt the same model architecture with the audio model of Voicebox (Fig. 2 of [18]) except that the frame-wise phoneme sequence is replaced into $\hat{y}$. Specifically, Transformer with U-Net style skip connection [18] is used as a backbone. The input to the mel spectrogram generator is $(1-m)\odot\hat{s}$, $\hat{y}$, the flow step $t$, and noisy speech $s_{t}$. $\hat{y}$ is first converted to character embedding sequence $\tilde{y}\in\mathbb{R}^{E\times T}$. Then, $(1-m)\odot\hat{s}$, $s_{t}$, $\tilde{y}$ are all stacked to form a tensor with a shape of $(2\cdot D+E)\times T$, followed by a linear layer to output a tensor with a shape of $D\times T$. Finally, an embedding representation, $\hat{t}\in\mathbb{R}^{D}$, of $t$ is appended to form the input tensor with a shape of $\mathbb{R}^{D\times(T+1)}$ to the Transformer. The Transformer is trained to output a vector field $v_{t}$ with the conditional flow-matching objective $\mathcal{L}^{\rm CFM}$.
@@ 2.5.1 Extension 1: Eliminating the need for transcription of audio prompts in inference
In certain application contexts, obtaining the transcription of the audio prompt can be challenging. To eliminate the requirement of the transcription of the audio prompt during inference, we introduce an extension, illustrated in Fig. 2. This extension, referred to as E2 TTS X1, assumes that we have access to the transcription of the masked region of the audio, which we use for $y$. During inference, the extended character sequence $\hat{y}^{\prime}$ is formed without $y^{\rm aud}$, namely,
$\hat{y}^{\prime}=(c^{\prime\prime}_{1},c^{\prime\prime}_{2},\ldots,c^{\prime\prime}_{M^{\rm text}},\underbrace{\langle F\rangle,\ldots,\langle F\rangle}_{\mathcal{T}\text{ times}}).$ (4)
@@
The transcription of the masked region of the training audio can be obtained in several ways. One method is to simply apply automatic speech recognition (ASR) to the masked region during training, which is straightforward but costly. In our experiment, we employed the Montreal Forced Aligner [28] to determine the start and end times of words within each training data sample. The masked region was determined in such a way that we ensured not to cut the word in the middle.
@@ 3.1 Training data
We utilized the Libriheavy dataset [30] to train our models. The Libriheavy dataset comprises 50,000 hours of read English speech from 6,736 speakers, accompanied by transcriptions that preserve case and punctuation marks. It is derived from the Librilight [31] dataset contains 60,000 hours of read English speech from over 7,000 speakers. For E2 TTS training, we used the case and punctuated transcription without any pre-processing. We also used a proprietary 200,000 hours of training data to investigate the scalability of the E2 TTS model.
@@ 3.2 Model configurations
We constructed our proposed E2 TTS models using a Transformer architecture. The architecture incorporated U-Net [23] style skip connections, 24 layers, 16 attention heads, an embedding dimension of 1024, a linear layer dimension of 4096. The character embedding vocabulary size was 399.555We used all characters and symbols that we found in the training data without filtering. The total number of parameters amounted to 335 million. We modeled the 100-dimensional log mel-filterbank features, extracted every 10.7 milliseconds from audio samples with a 24 kHz sampling rate. A BigVGAN [32]-based vocoder was employed to convert the log mel-filterbank features into waveforms. The masking length was randomly determined to be between 70% and 100% of the log mel-filterbank feature length during training. In addition, we randomly dropped all the conditioning information with a 20% probability for classifier-free guidance (CFG) [33]. All models were trained for 800,000 mini-batch updates with an effective mini-batch size of 307,200 audio frames. We utilized a linear decay learning rate schedule with a peak learning rate of $7.5\times 10^{-5}$ and incorporated a warm-up phase for the initial 20,000 updates. We discarded the training samples that exceeded 4,000 frames.
@@
In addition, we trained a regression-based duration model by following that of Voicebox [18]. It is based on a Transformer architecture, consisting of 8 layers, 8 attention heads, an embedding dimension of 512, a linear layer dimension of 2048. The training process involved 75,000 mini-batch updates with 120,000 frames. We used this duration model to estimate the target duration for a fair comparison with the Voicebox baseline. Note that we will also show that E2 TTS is robust for different target duration in Section 3.6.
@@ 3.3 Evaluation data and metrics
In order to assess our models, we utilized the test-clean subset of the LibriSpeech-PC dataset [35], which is an extension of LibriSpeech [36] that includes additional punctuation marks and casing. We specifically filtered the samples to retain only those with a duration between 4 and 10 seconds. Since LibriSpeech-PC lacks some of the utterances from LibriSpeech, the total number of samples was reduced to 1,132, sourced from 39 speakers. For the audio prompt, we extracted the last three seconds from a randomly sampled speech file from the same speaker for each evaluation sample. [^6] The test set used in our experiments can be accessed at: https://github.com/microsoft/e2tts-test-suite.
@@
Table 1: Objective results for LibriSpeech-PC test-clean evaluation set. WER is expressed in percentage. Our reproduction. LL, LH, and PP stand for Librilight, Libriheavy, and Proprietary, respectively.
@@
Table 3: Comparison between the basic E2 TTS and E2 TTS X1. While the basic E2 TTS requires the transcription of the audio prompt during inference, E2 TTS X1 does not require it. WER is expressed in percentage.
@@ 3.4 Main results
In our experiments, we conducted a comparison between our E2 TTS models and three other models: Voicebox [18], VALL-E [8], and NaturalSpeech 3 [15]. We utilized our own reimplementation of the Voicebox model, which was based on the same model configuration with E2 TTS except that the Vicebox model is trained with frame-wise phoneme alignment. During the inference, we used CFG with a guidance strength of 1.0 for both E2 TTS and Voicebox. We employed the midpoint ODE solver with a number of function evaluations of 32.
@@
Table 4: The WER (%) and SIM-o of E2 TTS X2 where a word is randomly replaced with the phoneme sequence during inference. Even when we replaced 50% of words into phoneme sequences, E2 TTS X2 worked reasonably well. This indicates that we can specify the pronunciation of a new term without retraining.
@@ 3.5.2 Evaluation of the extension 2
In this experiment, we trained the E2 TTS X2 models with a phoneme replacement rate of 15%. During inference, we randomly replaced words in the test-clean dataset with phoneme sequences, with a probability ranging from 0% to 50%.